# Edge Learning for 6G-enabled Internet of Things: A Comprehensive Survey of Vulnerabilities, Datasets, and Defenses

### v1

@@ VII-E Poisoning attacks
Poisoning attacks involve the manipulation of the training data to produce a biased or inaccurate model. Different types of poisoning attacks are presented, including federated poisoning attacks, intrusion poisoning attacks, label flipping attacks, poisonous label attacks, clean-label poisoning attacks, data poisoning attacks, and generative poisoning attacks [217, 160]. An attacker will penetrate a machine learning system and inject false or incorrect information into the database. Over time, as the algorithm learns from this falsified data, it will produce unwanted and potentially damaging outcomes [239, 240]. However, data poisoning attacks can be divided into two major classes: those attacking availability and others attacking integrity. Availability attacks are frequently not particularly advanced but extensive, inserting at least as many bad data items as they can into a database. Once an attack is successful, the machine learning algorithm will be totally incorrect. Attacks against the integrity of machine learning are more complicated and could be more dangerous.
@@ VII-E1 Federated poisoning attack
Federated learning uses machine learning in a decentralized mode that does not directly access the private data of clients. However, Federated learning suffers from many issues, such as adversarial machine learning-related security attacks, high communications costs between clients and the server, and high accuracy [157, 165, 160, 147, 10, 152, 161, 148]. To address these issues, Tabatabai et al. [171] present a federated learning algorithm based on evolutionary methods. The proposed clustering algorithm collects clients in many clusters, each with a selected model at random to exploit the capabilities of individual models. Then, the clusters are built in a repeated procedure with the worst cluster being dropped at every repetition process until there is only one cluster available. During all iterations, certain clients are kicked out of the clusters for corrupted data usage or poor overall performance. The winning clients are operated in the upcoming iteration. The cluster that remains with the surviving clients is then utilized for the best FL model training. The Fast Gradient Sign Method (FGSM) is adopted for producing adversarial examples in order to create a federated poisoning attack.
To mitigate data poisoning attacks, Doku and Rawat [173] proposed a security system in a federated learning mode. To generate the total error produced by a training dataset, the proposed system uses a Support Vector Machine. The performance evaluation with the IMDB review sentiment dataset shows that the proposed system achieved an accuracy score of $0.72$. Uprety and Rawat [156] proposed a technique for mitigating poisoning attacks that focuses on the reputation of the nodes involved in the process of training. The average reputation score of every client is computed with the method of probability distribution beta. The benchmark MNIST dataset is used in the performance evaluation, and the security analysis demonstrates that the classification model accuracy improved from $88$% to $93$% over $100$ rounds of communication after eliminating the attacker nodes by the aggregation servers. The generation of poison attacks in FL, as illustrated in Figure 10, can be defined by the following steps [175]:
@@ VII-E2 Intrusion poisoning attack
This type of attack consists of the application of poisoning attacks in the intrusion detection datasets Venkatesan et al. [213]. An intrusion detection system (IDS) is a network monitoring system that detects anomalous network behavior and issues alerts when anomalous activity is detected. An IDS is a computer program that analyzes a system or network for malicious activity or policy violations. However, in order to disturb the IDS systems, the attackers use intrusion poisoning attacks by implementing the poisoning attacks in the IDS dataset. The following datasets can be used for evaluating the performance of ML-based-IDS: Edge-IIoT dataset[56] and X-IIoTID dataset [241].
@@ VII-E3 Label flipping attack
This attack is conducted by flipping the labels of specific data samples from one class (the source class) to another (the target class), as shown in Figure 11. Jebreel et al. [174] introduced a new security system against the label-flipping attack in federated learning that extracts dynamically the target and source class potential gradients of local peer updates, performs an application of a clustering method on these gradients, and then evaluates the resulting groups to extract potential bad updates prior to the aggregation of the model. The following three data sets are used in the experiments: MNIST, CIFAR10, and IMDB. The results of three models CNN, ResNet18, and BiLSTM show that the proposed defense system can offer a lower attack success rate, higher source class accuracy, and lower test error. The following steps define the generation of label-flipping attacks in FL:
@@ VII-E4 Poisonous label attack
Compared to the poisonous attack under the white box requirement, the poisonous label attack is the black box, which is able to enhance the misclassification error under a more constrained and workable condition of the poisonous label attack. The poisonous label attack involves three parts: its capability, the attacker’s knowledge, and the attacker’s goal. The attackers’ goals are identified by three components: error specificity, attack specificity, and security violation. Liu et al. [155] propose the poisonous label attack, which injects fake images with the poisonous label in the training dataset than modifying directly the label of the images. The knowledge budget for the poisonous label attack is defined as:
@@ VII-E5 Clean-label poisoning attack
This is a class of poisoning attacks where the attacker has no control over the labeling procedure. In this particular threat model, the poison patterns are generated by inserting undetectable changes that will cause the model to misbehave in reaction to particular target inputs. Aghakhani et al. [151] propose a scalable clean-label poisoning attack, which enhances the current state-of-the-art attack success rate by 26.75% in transfer learning while improving the speed of attack by a factor of 12. This attack is based on injecting poisoned data into the training dataset in such a way that the model learns to classify the target input incorrectly. Based on the method that is based only on first-order information, Zheng et al. [150] proposed a clean-label data poisoning attack based on first-order information. The proposed attack is evaluated with the CIFAR-10 dataset on multiple network architectures, namely, ResNet, VGG, and ConvNet. The attack uses a Convex Polytope (CP) to solve an optimization problem, which involves minimizing the distance between the target feature vector and the poisoned samples set. The coefficients of the poisoned samples set are specified by $V^{i}_{j}$. The goal of the attack is to find a set of poisoned samples that can mislead the model when it encounters specific inputs. The CP adopted solves the following optimization problem:
@@
where $x^{j}_{p}$ is a poison samples set, ${\alpha}^{i}$ is the target feature vector, and $V^{i}_{j}$ specifies the $\ j$-th poison’s coefficient. The objective function is a convex combination of squared Euclidean distances, where each distance is weighted by a positive coefficient. The denominator of each term ensures that the objective function is normalized by the length of the target feature vector $\alpha^{i}(x_{t})$, which makes it invariant to the magnitude of the feature vector. The minimization problem seeks to find the value of $x_{t}$ that minimizes the objective function. Because the objective function is convex, any local minimum is also a global minimum. Therefore, any optimization algorithm that converges to a local minimum will also converge to the global minimum.
@@ VII-E6 Data poisoning attack
Data poisoning attacks against machine learning models have emerged as an influential research area of adversarial machine learning. This form of attack occurs during the training phase of machine learning models. The attacker attempts to infect the training data by inserting well-formed samples to impose a harmful model on the learner. Zhang et al. [242] proposed a data poisoning attack scheme, called IMF. The IMF scheme adopts the bi-level optimization problem to formulate the poisoning attack. Specifically, the upper-level problem estimates the strength of the actual fabricated poisoning sample, while the lower-level problem implements the model parameters based on these poisoning samples. The iterative learning algorithm is implemented by the IMF scheme to solve the bi-level optimization problem. The performance evaluation on two datasets, namely, the Stanford Dogs dataset and the Osteoporotic Fracture dataset, shows that the MFI scheme has the potential to reshape interpretations of the target samples successfully. However, Takiddin et al. [179] present a sequential ensemble detector against data poisoning attacks, which is based on feed-forward neural networks, gated recurrent units (GRUs), and a deep auto-encoder with attention (AEA).
@@ VII-E7 Generative poisoning attack
This type consists of deploying a generative adversarial network (GAN) model from the attacker’s side to impersonate other participants’ training dataset patterns, which can initiate the poisoning attack successfully under an assumption of a more feasible threat. Zhang et al. [153] propose a novel poisoning attack, named PoisonGAN, which can be applied in federated learning settings. Specifically, the PoisonGAN attack is based on generative adversarial networks as presented in Figure 10. Chen et al. [161] propose a novel adversarial network, named DeepPoison, which is based on one generator and two discriminators. More precisely, the generator extracts hidden features from the target class automatically and integrates them into benign training patterns. The one discriminator monitors the poisoning perturbation ratio. The second discriminator operates as a target model to witness the poisoning impacts. The idea of GAN is to build an adversarial game between a generator ${Gen}$ and a discriminator ${Disc}$, where ${Disc}$ is trained on the true samples and the generated samples simultaneously, causing the generator coupled ${Gen}$ to produce samples that are close to the true ones. The objective functions used by the GAN model are shown as follows:
@@ VII-E8 Over-the-air spectrum poisoning attack
This type of poisoning attack focuses on the spectrum sensing time period and handling the transmitter’s input data in the test and training stage. An adversary performs these attacks primarily to confuse the sender into making poor transmission decisions (i.e., an evasion attack) or to tamper with the sender’s recycling process (i.e., a causal attack). To carry out an over-the-air spectrum poisoning attack, an attacker needs to have knowledge of the transmitter’s sensing parameters and the communication protocol being used. Once the attacker has this information, he can craft adversarial examples that exploit the vulnerabilities of the transmitter’s decision-making process. Sagduyu et al. [214] present over-the-air spectrum sensing poisoning attacks based on the application of adversarial machine learning. Benchmarking shows that over-the-air spectrum sensing poisoning attacks are superior to conventional jamming attacks and substantially decrease the transmitter throughput. Therefore, over-the-air spectrum poisoning attacks can be more effective than traditional jamming attacks, which simply flood the communication channel with noise to disrupt communication. Additionally, these attacks can significantly reduce the throughput of the transmitter, which can have significant consequences in applications such as wireless networks and IoT devices.
TABLE XV: Defense methods against machine learning vulnerabilities.
@@ VII-F Sybil attacks
Sybil attacks involve the creation of multiple fake identities to manipulate the training process of a federated learning system. Different types of Sybil attacks are presented, including Sybil-based poisoning attacks and Sybil-based collusion attacks [147]. A Sybil attack involves the use of a single node inside the targeted network (or system) to simultaneously generate and operate multiple active false entities. The primary objective of these attacks is to gain the majority of influence within the targeted system in order to facilitate its manipulation.
@@ VII-F1 Sybil-based poisoning attack
Sybils conduct federated learning poisoning attacks by delivering status updates that lead the distributed model to a poisoned common target. Based on contribution similarity, Fung et al. [147] propose a new defense system against Sybil-based poisoning attacks. The proposed system adapts the learning rate of clients based on contribution similarity.
@@ VII-F2 Sybil-based collusion attack
Xiao et al. [166] propose the Sybil-based collusion attack, in which the attacker employs the label flipping scheme to train the poison data locally and collide with other poisoned patterns. The malicious adversary, meanwhile, will virtually implement several Sybil nodes in the network, so that the server chooses the collusion pattern to aggregate with a higher probability, constructing a poisoning pattern globally. Both CIFAR-10 and Fashion-MNIST are used as benchmark datasets to evaluate the attack performance in federated learning settings with convolutional Neural Networks (CNN). The experiment analysis shows that the proposed attack can achieve a more substantial attack effect.
@@ VII-G Byzantine attacks
These attacks consist of Byzantine failures in federated learning, which can exploit the parameters of the local model on the affected devices during the learning phase. Specifically, Byzantine attacks in federated learning involve malicious participants that deliberately manipulate the local model to corrupt the training process. These attacks can severely compromise the integrity and accuracy of the final global model. The participants that perform Byzantine attacks are known as Byzantine faulty or malicious participants. These participants can exhibit arbitrary behavior and can send incorrect or intentionally corrupted updates to the central server, leading to the poisoning of the global model. The attacker’s knowledge in Byzantine attacks can be classified into three levels: No knowledge, Partial knowledge, and Complete knowledge. In the case of no knowledge, the attacker has no specific knowledge of the distributed machine learning model, its architecture, or the data it’s working with. An attacker with partial knowledge has some information about the distributed machine learning model or its architecture, but not complete details. In the scenario of complete knowledge, the attacker has access to all information related to the distributed machine learning system, including its architecture, model parameters, training data, and even updates from other nodes.

### v2

@@ VII-E Poisoning attacks
Poisoning attacks involve manipulating the training data to produce a biased or inaccurate model. Different types of poisoning attacks are presented, including federated poisoning attacks, intrusion poisoning attacks, label flipping attacks, poisonous label attacks, clean-label poisoning attacks, data poisoning attacks, and generative poisoning attacks [244, 181]. An attacker will penetrate a machine learning system and inject false or incorrect information into the database. Over time, as the algorithm learns from this falsified data, it will produce unwanted and potentially damaging outcomes [266, 267]. However, data poisoning attacks can be divided into two major classes: those attacking availability and others attacking integrity. Availability attacks are frequently not particularly advanced but extensive, inserting at least as many bad data items as possible into a database. Once an attack is successful, the machine learning algorithm will be incorrect. Attacks against the integrity of machine learning are more complicated and could be more dangerous.
@@ VII-E1 Federated poisoning attack
Federated learning uses machine learning in a decentralized mode that does not directly access clients’ private data. However, Federated learning suffers from many issues, such as adversarial machine learning-related security attacks, high communications costs between clients and the server, and high accuracy [178, 186, 181, 168, 13, 173, 182, 169]. Federated poisoning attacks can manipulate the aggregated model in 6G by introducing malicious updates from compromised devices. This could degrade the performance and reliability of 6G network services that rely on machine learning. As 6G networks are anticipated to be more AI-driven, such attacks may pose significant threats to network security and functionality.To address these issues, Tabatabai et al. [191] present a federated learning algorithm based on evolutionary methods. The proposed clustering algorithm collects clients in many clusters, each with a selected model at random to exploit the capabilities of individual models. Then, the clusters are built in a repeated procedure with the worst cluster being dropped at every repetition process until only one cluster is available. Specific clients are kicked out of the clusters for corrupted data usage or poor overall performance during all iterations. The winning clients are operated in the upcoming iteration. The cluster that remains with the surviving clients is then utilized for the best FL model training. The Fast Gradient Sign Method (FGSM) is adopted for producing adversarial examples to create a federated poisoning attack.
To mitigate data poisoning attacks, Doku and Rawat [193] proposed a security system in a federated learning mode. The proposed system uses a Support Vector Machine to generate the total error a training dataset produces. The performance evaluation with the IMDB review sentiment dataset shows that the proposed system achieved an accuracy score of $0.72$. Uprety and Rawat [177] proposed a technique for mitigating poisoning attacks that focuses on the nodes’ reputation in the training process. The average reputation score of every client is computed with the method of probability distribution beta. The benchmark MNIST dataset is used in the performance evaluation, and the security analysis demonstrates that the classification model accuracy improved from $88$% to $93$% over $100$ rounds of communication after eliminating the attacker nodes by the aggregation servers. The generation of poison attacks in FL, as illustrated in Figure 12, can be defined by the following steps [195]:
@@ VII-E2 Intrusion poisoning attack
Intrusion poisoning attacks can compromise 6G networks by feeding misleading data into the network’s self-learning systems, leading to suboptimal decisions and vulnerabilities. Given 6G’s reliance on AI-driven functionalities and ultra-reliable low-latency communication, malicious data injection can disrupt real-time operations. This attack can erode trust in the network’s intelligence, undermining its promise of enhanced connectivity and smarter applications. Specifically, the intrusion poisoning attack consists of applying poisoning attacks in the intrusion detection datasets Venkatesan et al. [233]. An intrusion detection system (IDS) is a network monitoring system that detects anomalous network behavior and alerts when anomalous activity is detected. An IDS is a computer program that analyzes a system or network for malicious activity or policy violations. However, to disturb the IDS systems, the attackers use intrusion poisoning attacks by implementing the poisoning attacks in the IDS dataset. The following datasets can be used for evaluating the performance of ML-based-IDS: Edge-IIoT dataset[123] and X-IIoTID dataset [268].
@@ VII-E3 Label flipping attack
Label flipping attacks involve maliciously altering the data labels to mislead training processes. In the context of 6G, which promises enhanced machine learning and AI integrations, a label-flipping attack could compromise the integrity of AI-driven network functions. Consequently, this could degrade network performance, compromise user experience, and potentially introduce vulnerabilities to other integrated systems. Therefore, this attack is conducted by flipping the labels of specific data samples from one class (the source class) to another (the target class), as shown in Figure 13. Jebreel et al. [194] introduced a new security system against the label-flipping attack in federated learning that dynamically extracts the target and source class potential gradients of local peer updates, performs an application of a clustering method on these gradients, and then evaluates the resulting groups to extract potential bad updates before the aggregation of the model. The following three data sets are used in the experiments: MNIST, CIFAR10, and IMDB. The results of three models CNN, ResNet18, and BiLSTM show that the proposed defense system can offer a lower attack success rate, higher source class accuracy, and lower test error. The following steps define the generation of label-flipping attacks in FL:
@@ VII-E4 Poisonous label attack
Compared to the poisonous attack under the white box requirement, the poisonous label attack is the blue box, which can enhance the misclassification error under a more constrained and workable condition of the poisonous label attack. A Poisonous label attack on 6G could manipulate training data, skewing the model’s understanding and behavior. As 6G heavily relies on AI-driven systems for optimization and decision-making, such attacks could compromise network performance and security. Malicious actors could exploit this vulnerability, causing communication disruptions or unauthorized data access. The poisonous label attack involves its capability, the attacker’s knowledge, and the attacker’s goal. The attackers’ goals are identified by three components: error specificity, attack specificity, and security violation. Liu et al. [176] propose the poisonous label attack, which injects fake images with the poisonous label in the training dataset than directly modifying the label of the images. The knowledge budget for the poisonous label attack is defined as:
@@ VII-E5 Clean-label poisoning attack
This is a class of poisoning attacks where the attacker has no control over the labeling procedure. In this particular threat model, the poison patterns are generated by inserting undetectable changes that will cause the model to misbehave in reaction to particular target inputs. Clean-label poisoning attacks introduce subtly modified, malicious samples into the training data of machine learning models. In the context of 6G, which heavily relies on AI for network optimization and decision-making, such attacks could degrade network performance, compromise security protocols, and disrupt intelligent service provisions. The ubiquity and complexity of 6G networks amplify the potential impact and challenges of defending against these sophisticated threats. Aghakhani et al. [172] propose a scalable clean-label poisoning attack, which enhances the current state-of-the-art attack success rate by 26.75% in transfer learning while improving the speed of attack by a factor of 12. This attack is based on injecting poisoned data into the training dataset so that the model learns to classify the target input incorrectly. Based on the method that is based only on first-order information, Zheng et al. [171] proposed a clean-label data poisoning attack based on first-order information. The proposed attack is evaluated with the CIFAR-10 dataset on multiple network architectures: ResNet, VGG, and ConvNet. The attack uses a Convex Polytope (CP) to solve an optimization problem involving minimizing the distance between the target feature vector and the poisoned samples set. The coefficients of the poisoned samples set are specified by $V^{i}_{j}$. The attack aims to find a set of poisoned samples that can mislead the model when it encounters specific inputs. The CP adopted solves the following optimization problem:
@@
where $x^{j}_{p}$ is a poison samples set, ${\alpha}^{i}$ is the target feature vector, and $V^{i}_{j}$ specifies the $\ j$-th poison’s coefficient. The objective function is a convex combination of squared Euclidean distances, where a positive coefficient weights each distance. The denominator of each term ensures that the objective function is normalized by the length of the target feature vector $\alpha^{i}(x_{t})$, which makes it invariant to the magnitude of the feature vector. The minimization problem seeks to find the value of $x_{t}$ that minimizes the objective function. Because the objective function is convex, any local minimum is also a global minimum. Therefore, any optimization algorithm that converges to a local minimum will also converge to the global minimum.
@@ VII-E6 Data poisoning attack
Data poisoning attacks against machine learning models have emerged as an influential adversarial machine learning research area. This form of attack occurs during the training phase of machine learning models. Data poisoning attacks in 6G can introduce maliciously crafted data into the network’s training datasets, undermining the AI-driven functions vital for 6G optimizations. As 6G relies heavily on machine learning for various processes, tainted data can lead to misinformed network decisions and potential service disruptions. The attacker attempts to infect the training data by inserting well-formed samples to impose a harmful model on the learner. Zhang et al. [269] proposed a data poisoning attack scheme called IMF. The IMF scheme adopts the bi-level optimization problem to formulate the poisoning attack. Specifically, the upper-level problem estimates the strength of the actual fabricated poisoning sample, while the lower-level problem implements the model parameters based on these poisoning samples. The iterative learning algorithm is implemented by the IMF scheme to solve the bi-level optimization problem. The performance evaluation on two datasets, namely, the Stanford Dogs dataset and the Osteoporotic Fracture dataset, shows that the MFI scheme has the potential to reshape interpretations of the target samples successfully. However, Takiddin et al. [199] present a sequential ensemble detector against data poisoning attacks, which is based on feed-forward neural networks, gated recurrent units (GRUs), and a deep auto-encoder with attention (AEA).
@@ VII-E7 Generative poisoning attack
This type consists of deploying a generative adversarial network (GAN) model from the attacker’s side to impersonate other participants’ training dataset patterns, which can initiate the poisoning attack successfully under an assumption of a more feasible threat. A GAN-based poisoning attack in a 6G environment involves adversarial manipulation of training datasets, misleading network algorithms into making incorrect decisions. Zhang et al. [174] propose a novel poisoning attack, PoisonGAN, which can be applied in federated learning settings. Specifically, the PoisonGAN attack is based on generative adversarial networks as presented in Figure 12. Chen et al. [182] propose a novel adversarial network, DeepPoison, based on one generator and two discriminators. More precisely, the generator automatically extracts hidden features from the target class and integrates them into benign training patterns. The one discriminator monitors the poisoning perturbation ratio. The second discriminator operates as a target model to witness the poisoning impacts. The idea of GAN is to build an adversarial game between a generator ${Gen}$ and a discriminator ${Disc}$, where ${Disc}$ is trained on the actual samples and the generated samples simultaneously, causing the generator coupled ${Gen}$ to produce samples that are close to the true ones. The objective functions used by the GAN model are shown as follows:
@@ VII-E8 Over-the-air spectrum poisoning attack
Over-the-air spectrum poisoning attacks in 6G can maliciously introduce interference in the radio frequency spectrum, leading to degraded network performance. By exploiting the intelligent and adaptive nature of 6G’s signal processing, adversaries can inject, amplify, or manipulate spectrum data. This disrupts the AI-driven operations in 6G networks, potentially leading to unreliable communication, misallocated resources, or complete service outages. This poisoning attack focuses on the spectrum sensing period and handling the transmitter’s input data in the test and training stage. An adversary performs these attacks primarily to confuse the sender into making poor transmission decisions (i.e., an evasion attack) or to tamper with the sender’s recycling process (i.e., a causal attack). To carry out an over-the-air spectrum poisoning attack, an attacker must know the transmitter’s sensing parameters and the communication protocol used. Once the attacker has this information, he can craft adversarial examples that exploit the vulnerabilities of the transmitter’s decision-making process. Sagduyu et al. [234] present over-the-air spectrum sensing poisoning attacks based on the application of adversarial machine learning. Benchmarking shows that over-the-air spectrum sensing poisoning attacks are superior to conventional jamming attacks and substantially decrease the transmitter throughput. Therefore, over-the-air spectrum poisoning attacks can be more effective than traditional jamming attacks, which flood the communication channel with noise to disrupt communication. Additionally, these attacks can significantly reduce the throughput of the transmitter, which can have significant consequences in applications such as wireless networks and IoT devices.
TABLE XVI: Defense methods against machine learning vulnerabilities.
@@ VII-F Sybil attacks
Sybil attacks involve the creation of multiple fake identities to manipulate the training process of a federated learning system. Different types of Sybil attacks are presented, including Sybil-based poisoning attacks and Sybil-based collusion attacks [168]. A Sybil attack involves using a single node inside the targeted network (or system) to generate and operate multiple active false entities simultaneously. The primary objective of these attacks is to gain the majority of influence within the targeted system to facilitate its manipulation.
@@ VII-F1 Sybil-based poisoning attack
A Sybil-based poisoning attack in 6G networks involves malicious actors creating multiple fake identities to influence the network’s decision-making processes. By overwhelming the network with deceptive information from these fake nodes, the attack can degrade the quality of service, data integrity, and trustworthiness. This can lead to compromised network performance, erroneous decisions, and reduced security in the 6G ecosystem. Sybils conduct federated learning poisoning attacks by delivering status updates that lead the distributed model to a poisoned common target. Based on contribution similarity, Fung et al. [168] propose a new defense system against Sybil-based poisoning attacks. The proposed system adapts the learning rate of clients based on contribution similarity.
@@ VII-F2 Sybil-based collusion attack
In 6G networks prioritizing decentralized, AI-driven infrastructures and massive device connectivity, Sybil-based collusion attacks can introduce numerous malicious identities, misleading the system’s decision-making processes. By artificially amplifying specific metrics or information, attackers can disrupt network services, degrade trustworthiness, and compromise AI-driven functionalities. Xiao et al. [187] propose the Sybil-based collusion attack, in which the attacker employs the label flipping scheme to train the poison data locally and collide with other poisoned patterns. Meanwhile, The malicious adversary will virtually implement several Sybil nodes in the network so that the server chooses the collusion pattern to aggregate with a higher probability, constructing a global poisoning pattern. CIFAR-10 and Fashion-MNIST are benchmark datasets to evaluate attack performance in federated learning settings with convolutional Neural Networks (CNN). The experiment analysis shows that the proposed attack can achieve a more substantial attack effect.
@@ VII-G Byzantine attacks
These attacks consist of Byzantine failures in federated learning, which can exploit the parameters of the local model on the affected devices during the learning phase. Specifically, Byzantine attacks in federated learning involve malicious participants that deliberately manipulate the local model to corrupt the training process. These attacks can severely compromise the integrity and accuracy of the final global model. The participants performing Byzantine attacks are known as Byzantine faulty or malicious. These participants can exhibit arbitrary behavior and send incorrect or intentionally corrupted updates to the central server, leading to the poisoning of the global model. The attacker’s knowledge of Byzantine attacks can be classified into No, Partial, and Complete. Without knowledge, the attacker has no specific knowledge of the distributed machine learning model, its architecture, or the data it’s working with. An attacker with partial knowledge has some information about the distributed machine learning model or its architecture but not complete details. In the scenario of complete knowledge, the attacker can access all information related to the distributed machine learning system, including its architecture, model parameters, training data, and even updates from other nodes.